OpenStudy (anonymous):

plz help will fan and medal if correct How many solutions does the equation have? 6-a=1/3(9a-18)

OpenStudy (anonymous):



OpenStudy (anonymous):

plz hurry anyone plz help

OpenStudy (anonymous):

you can simplify the question by distributing over the parentheses. multiplying by 1/3 is like dividing by 3, as 3 is the reciprocal of 1/3. so to simplify, 6 - a = 3a - 6. we then can put like terms on the same side, so 6 - a + a = 3a + a - 6, 6 = 4a - 6; 6 + 6 = 4a - 6 + 6, 12 = 4a. Then by dividing, 12/4 = 4a/4, 3 = a.

OpenStudy (anonymous):

so a = 3?

OpenStudy (anonymous):

yes. This would mean that the equation has only one solution, as only one number can be put into a for the problem to be correct

OpenStudy (anonymous):

oh wait I forgot to add somethin' A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

oh ok thx can u help with 3 more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thx close and open a new one?

OpenStudy (anonymous):

okay